If Kc = 0.440 at 40.°C and Kc = 0.515 at 90.°C, what is ΔH° for the reaction?

X <---> Y
Answer #1

Using,

ln(k2/k1) = dHo/R[1/T1-1/T2]

where,

k1 = 0.440

k2 = 0.515

T1 = 40 oC = 40 + 273 = 313 K

T2 = 90 oC = 90 + 273 = 363 K

R = 8.314 J/K.mol

Feed values,

ln(0.515/0.440) = dHo/8.314[1/313-1/363]

dH = 2973.54 J = 3.0 kJ

So deltaHo (dHo) for the reaction is 3 kJ